OpenStudy (anonymous):

will fan and medal !! a sequence is defined recursively using the formula f(n+1)=-0.5f(n). if the first term of the sequence is 120, what is f(5) ?

OpenStudy (sidsiddhartha):

start with putting n=1

OpenStudy (sidsiddhartha):

it will gove u f(2)

OpenStudy (sidsiddhartha):

*give

OpenStudy (anonymous):

I don't know how to do this at all , im just lookin for the answer so I can pass this test lol

OpenStudy (sidsiddhartha):

n=1 gives-\[f(1+1)=-0.5f(1)\\f(2)=-0.5*120\] geting this?

OpenStudy (anonymous):

noo lol

OpenStudy (sidsiddhartha):

so f(2)=-60 ok?

OpenStudy (anonymous):

lol can I have just the answer ?

OpenStudy (sidsiddhartha):

keep doing this next put n=2 \[f(2+1)=-0.5*f(2)=-0.5*-60=30\\f(3)=30\]

OpenStudy (sidsiddhartha):

next, put n=3\[f(4)=-0.5f*(3)=-0.5*30=-15\] and finally put n=4 \[f(5)=-0.5*f(4)=7.5\] so f(5)=7.5

OpenStudy (sidsiddhartha):

thats ur answer

OpenStudy (anonymous):

thank youuuu !

OpenStudy (sidsiddhartha):

yw!!